OpenStudy (littlexmonster):

A prism is made of cubes which measure 1/3 of an inch on one side. What is the volume?

OpenStudy (confluxepic):

Volume is side X side X side.

OpenStudy (confluxepic):

\[V=s \times s \times s\]

OpenStudy (confluxepic):

S=? What does the question say?

OpenStudy (confluxepic):

@LittlexMonster

OpenStudy (littlexmonster):

What I posted was the complete question

OpenStudy (littlexmonster):

That's all it gave me

OpenStudy (visarika23):

S=1/3

OpenStudy (confluxepic):

Yes. But use information from the question and plug it into the equation.

OpenStudy (confluxepic):

@LittlexMonster What is the side measure?

OpenStudy (littlexmonster):

Answers it gave me 7/3 cubic in , 12cubic in , 4/9 cubic in, 4 cubic in

OpenStudy (confluxepic):

How did you get that?

OpenStudy (confluxepic):

Take the side and cube it. s=1/3 V=s^3 =(1/3)^3

OpenStudy (littlexmonster):

It is multiple choice those were the choices

OpenStudy (littlexmonster):

*multiple question

OpenStudy (confluxepic):

Do you have an image of the prism that contains the cube?

OpenStudy (littlexmonster):

Yes

OpenStudy (welshfella):

can you show it?

OpenStudy (confluxepic):

Can you post it? Otherwise I won't be able to solve the question.

OpenStudy (welshfella):

The volume of the prism will depend on the number of cubes in it.

OpenStudy (confluxepic):

\[(\frac{ 1 }{ 3 })^3=\frac{ 1^3 }{ 3^3 }=\frac{ 1 }{ 27 }\]

OpenStudy (confluxepic):

Then multiply that by the number of cubes in the prism to get your answer.